Q.8

Attempts: Score: /3 8. Problems and Applications Q8 GDP Year 2011 2012 2013 2014 2015 (Billions of current dollars) 15,517.9 16,155.3 16,663.2 17,348.1 17,942.9 (Billions of 2009 dollars) 15,020.6 15,354.6 15,583.3 15,961.7 16,345.0 Source: National Economic Accounts.U.S. Bureau of Economic Analysis, The percentage change in nominal GDp from 2011 to 2012 was. The percentage change in real GDP from 2011 to 2012 wasY True or False: The percentage change in real GDP from 2011 to 2012 was higher than the percentage change in real GDP from 2012 to 2013. O True False
Answer #1

GDP at current billions of dollars are nominal GDP while GDP at billions of 2009 dollars are real GDP.

Nominal GDP in 2011 = 15,517.9

Nominal GDP in 2012 =16,155.3

Change in nominal GDP = Nominal GDP in 2012 - Nominal GDP in 2011

= 16,155.3 - 15,517.9

= 637.4

Percentage change in nominal GDP between 2011 and 2012

= (Change in nominal GDP/ Nominal GDP in 2011)*100

=(637.4/15,517.9)*100

=4.107%

Real GDP in 2011 = 15,020.6

Real GDP in 2012 =15,354.6

Change in real GDP = Real GDP in 2012 - Real GDP in 2011

= 15,354.6 - 15,020.6

= 334

Percentage change in Real GDP between 2011 and 2012

= (Change in Real GDP/ Real GDP in 2011)*100

=(334/15,020.6)*100

=2.22%

Real GDP in 2012 =15,354.6

Real GDP in 2013 = 15,583.3

Change in real GDP = Real GDP in 2013 - Real GDP in 2012

=15,583.3- 15,354.6

= 228.7

Percentage change in Real GDP between 2012 and 2013

= (Change in Real GDP/ Real GDP in 2012)*100

=(228.7/15,354.6 )*100

   =1.48%

Percentage change in Real GDP between 2012 and 2013 is 1.48%

Percentage change in Real GDP between 2011 and 2012 is 2.22%

As it can be seen that percentage change in real GDP between 2011 and 2012 is higher than the percentage change in real GDP between 2012 and 2013.

Therefore the given statement is true.